### v1

# Towards Source Mapping for Zero-Knowledge Smart Contracts: Design and Preliminary Evaluation

Abstract.
Debugging and auditing zero-knowledge-compatible smart contracts remains a significant challenge due to the lack of source mapping in compilers such as zkSolc. In this work, we present a preliminary source mapping framework that establishes traceability between Solidity source code, LLVM IR, and zkEVM bytecode within the zkSolc compilation pipeline. Our approach addresses the traceability challenges introduced by non-linear transformations and proof-friendly optimizations in zero-knowledge compilation.
To improve the reliability of mappings, we incorporate lightweight consistency checks based on static analysis and structural validation. We evaluate the framework on a dataset of 50 benchmark contracts and 500 real-world zkSync contracts, observing a mapping accuracy of approximately 97.2%  for standard Solidity constructs. Expected limitations arise in complex scenarios such as inline assembly and deep inheritance hierarchies. The measured compilation overhead remains modest, at approximately 8.6% .
Our initial results suggest that source mapping support in zero-knowledge compilation pipelines is feasible and can benefit debugging, auditing, and development workflows. We hope that this work serves as a foundation for further research and tool development aimed at improving developer experience in zk-Rollup environments.
Zero-Knowledge Proofs, zk-Rollups, zkSolc, Smart Contracts, Source Mapping, Static Analysis, Compilation Pipeline
conference: The First Workshop on EXPlainable and REliable Software Systems (EXPRESS 2025); June 2025; Trondheim, Norwayccs: Software and its engineering Compilersccs: Security and privacy Formal security modelsccs: Theory of computation Program analysis
1. Introduction
Blockchain technology has become a cornerstone of decentralized digital asset systems, enabling transparent, secure, and efficient applications (Kushwaha et al., 2022). Among the solutions addressing blockchain scalability challenges, zk-Rollups have emerged as a promising technology. They allow Ethereum’s Layer 2 ecosystem to process large transaction volumes while preserving strong security guarantees (Xu et al., 2022). A key enabler of zk-Rollups is zkSolc, a specialized compiler that extends the Solidity compiler to produce zero-knowledge-compatible bytecode, commonly referred to as zkEVM bytecode.
Source mapping refers to the process of correlating compiled bytecode instructions with their corresponding high-level source code locations. This mapping is essential for debugging, auditing, and symbolic execution, as it enables developers and auditors to trace execution behavior back to source code statements.
While source mapping is a standard feature in Solidity compilers such as solc, it is currently absent in zkSolc. This omission introduces practical challenges for developers working on zero-knowledge-based smart contracts. Specifically, the zkSolc compilation pipeline involves multiple transformation stages, including translation to LLVM IR and subsequent generation of zkEVM bytecode. These stages introduce non-linear transformations, such as instruction reordering, arithmetic simplification, and control flow restructuring, which break the direct correspondence between source code and bytecode offsets that traditional source mapping techniques rely on.
In this work, we propose a preliminary source mapping framework for zkSolc. Our objective is to improve traceability across compilation stages and facilitate debugging and auditing in zk-Rollup environments.
1.1. Contributions
This paper makes the following contributions:
  • We propose a source mapping framework for zkSolc that links Solidity source code, LLVM IR, and zkEVM bytecode, accounting for transformations specific to zero-knowledge compilation.
  • We introduce a lightweight consistency validation process to assess the structural integrity of the generated mappings.
  • We report preliminary evaluation results on benchmark and real-world smart contracts, analyzing mapping accuracy and compilation overhead.
The remainder of this paper is organized as follows. Section 2 presents a motivating example. Section 3 introduces key concepts and defines the problem. Section 4 describes the proposed framework. Section 5 reports evaluation results. Section 6 discusses limitations. Section 7 reviews related work, and Section 8 concludes the paper.
2. Motivating Example
To illustrate the importance of source mapping in zkSolc and the practical challenges that arise in its absence, we present a debugging scenario involving a privacy-preserving voting contract. This contract uses zero-knowledge proofs (zk-SNARKs) to protect voter anonymity while ensuring vote integrity. A simplified Solidity implementation is shown below:
1contract ZKVoting {
2 mapping(address => bool) public hasVoted;
3
4 function submitVote(bytes memory zkProof) external {
5 require(verifyZKProof(zkProof), "Invalid proof");
6 require(!hasVoted[msg.sender], "Already voted");
7
8 hasVoted[msg.sender] = true;
9 }
10
11 function verifyZKProof(bytes memory zkProof) internal pure returns (bool) {
12 // Placeholder for zero-knowledge proof verification logic
13 return true; // Simplified for illustration
14 }
15}
2.1. Debugging Challenges Without Source Mapping
In practice, debugging smart contracts compiled with zkSolc is challenging due to the absence of source mapping. Consider the following scenario: A user submits a vote, but the transaction fails at the require statement in the submitVote function. The error message "Invalid proof" indicates that verifyZKProof returned false. Without source mapping, identifying the root cause is difficult and may stem from:
  1. (1)
    A bug in the Solidity implementation of verifyZKProof.
  2. (2)
    An unintended transformation during the LLVM IR compilation stage.
  3. (3)
    A miscompilation or instruction reordering in the LLVM IR to zkEVM bytecode translation.
Without a mapping mechanism, developers must manually inspect low-level representations. For example, the corresponding LLVM IR may appear as:
1; Check the result of verifyZKProof
2%3 = call i1 @verifyZKProof(%bytes* %zkProof)
3br i1 %3, label %4, label %5
While this snippet shows a conditional branch based on verifyZKProof, it provides no direct connection to the original Solidity source. Similarly, the associated zkEVM bytecode:
10x03 PUSH1 0x40
20x04 MSTORE
30x05 CALLDATALOAD
lacks context or information about its origin in the Solidity logic. This absence of traceability complicates debugging and increases the manual effort required to locate errors.
2.2. Implications for Security Audits and Optimization
Beyond debugging, source mapping plays an essential role in security auditing and performance analysis. Auditors inspecting for vulnerabilities such as reentrancy or access control flaws often analyze low-level bytecode to assess contract behavior. Without source mapping, auditors must manually correlate bytecode instructions with Solidity source code, increasing the risk of misinterpretation.
Similarly, developers seeking to optimize gas costs require a clear mapping between expensive bytecode instructions and their high-level Solidity counterparts. For example, a sequence of bytecode instructions such as:
10x10 CALLDATASIZE
20x11 ISZERO
30x12 PUSH2 0x0040
40x13 JUMPI
could correspond to multiple Solidity statements, making it difficult to attribute gas consumption without source-level context.
2.3. Challenges in Implementing Source Mapping in zkSolc
While source mapping is a standard feature in traditional EVM compilers, implementing it in zkSolc presents unique challenges. The zkSolc compilation pipeline introduces multiple transformation stages, including lowering Solidity to LLVM IR, applying proof-friendly optimizations, and generating zkEVM-compatible bytecode. These transformations disrupt the direct correspondence between source code and compiled output.
For instance, a simple bitwise operation in Solidity:
1function checkBit(uint256 input) public pure returns (bool) {
2 return (input & 1) == 1;
3}
may be transformed during compilation into arithmetic constraints used in zk-SNARK proof generation:
1Constraint 1: tmp1 = input % 2
2Constraint 2: tmp2 = tmp1 == 1
These transformations improve proof efficiency but obscure the original source semantics, making conventional source mapping infeasible.
Additionally, LLVM optimizations such as inlining, constant folding, and dead code elimination further alter or eliminate source-level constructs. Witness generation in zero-knowledge proofs imposes bit-level precision requirements that introduce further rewrites, complicating traceability.
Performance overhead is another consideration. Including detailed source mapping increases the size of compiled artifacts and introduces additional processing steps during compilation, potentially impacting deployment costs and throughput.
Despite these challenges, we believe that source mapping support in zkSolc is an important direction. Improved traceability can assist developers and auditors in understanding contract behavior, identifying errors, and analyzing performance. This work represents an initial attempt to address this gap by proposing a source mapping framework tailored to the zkSolc compilation pipeline.
3. Preliminaries
This section introduces key background concepts relevant to our work, including zk-Rollups, the zkSolc compiler, and the challenges associated with source mapping in zero-knowledge compilation pipelines.
3.1. Blockchain Scaling and zk-Rollups
Blockchain scalability remains a central challenge. On Ethereum, limited transaction throughput and high gas fees have driven the development of Layer 2 scaling solutions that execute transactions off-chain while periodically settling state changes on-chain (Rebello et al., 2024; Rao et al., 2024). Among these, zk-Rollups leverage zero-knowledge proofs to validate the correctness of batched transactions without requiring full re-execution on the base layer.
In a typical zk-Rollup system, transactions are aggregated off-chain, and a succinct zk-SNARK proof is generated to attest to the correctness of the batch. This proof is submitted to an Ethereum smart contract, which verifies it efficiently. Unlike optimistic rollups that rely on fraud proofs and challenge periods, zk-Rollups enforce correctness at the time of state transition, providing immediate finality and strong security guarantees.
3.2. zkSolc: The Zero-Knowledge Solidity Compiler
zkSolc is a specialized compiler that extends the standard Solidity compiler to generate bytecode compatible with zero-knowledge proof systems. Specifically, it produces zkEVM bytecode, a variant of EVM bytecode optimized for zero-knowledge execution.
Unlike solc, which directly compiles Solidity source code to EVM bytecode, zkSolc introduces an intermediate compilation stage based on LLVM IR. This additional stage enables advanced optimizations tailored to zk-SNARK constraint generation but introduces substantial transformations that complicate traceability between source code and compiled output.
Figure 1. zkSolc compilation pipeline (Labs, 2025a). Yul is an intermediate language used in Solidity compilation. Vyper and LLL are alternative smart contract languages supported by EVM toolchains but are outside the scope of this work.
As illustrated in Figure 1, the zkSolc compilation pipeline consists of the following stages:
  1. (1)
    Parsing Solidity source code into an abstract syntax tree (AST).
  2. (2)
    Lowering the AST to a high-level intermediate representation (Yul).
  3. (3)
    Translating Yul to LLVM IR in static single assignment (SSA) form.
  4. (4)
    Applying LLVM optimizations to simplify arithmetic operations, restructure control flow, and minimize circuit constraints.
  5. (5)
    Generating zkEVM-compatible bytecode suitable for zero-knowledge execution.
These transformations improve proof efficiency but significantly alter instruction structure and semantics compared to the original Solidity source code. As a result, conventional offset-based source mapping techniques used in solc are not directly applicable.
3.3. Source Mapping in Solidity Compilers
Source mapping is a standard feature in traditional smart contract compilers, providing a correspondence between compiled bytecode instructions and their source-level origins. In conventional Solidity compilers such as solc, source maps are generated to facilitate debugging, symbolic execution, and security auditing. These maps enable developers to trace execution behavior back to Solidity source lines, supporting tools such as Remix IDE (Project, 2025b), Hardhat Debugger (hardhat), and Truffle (Suite, 2025), as well as formal analysis engines (of Bits, 2025c; ConsenSys, 2025).
A typical source map records instruction offsets, source file indices, and control flow metadata to help reconstruct execution traces and detect vulnerabilities such as reentrancy or access control violations.
3.4. Challenges in Source Mapping for Zero-Knowledge Compilation
Implementing source mapping in zkSolc introduces several challenges absent in conventional EVM compilation:
  • The introduction of LLVM IR as an intermediate representation changes instruction layout and semantics.
  • LLVM-based optimizations reorder, inline, and eliminate Solidity constructs to improve proof efficiency.
  • Zero-knowledge-specific rewrites further restructure control flow and data dependencies to reduce circuit complexity.
These transformations disrupt the direct relationship between Solidity code and compiled zkEVM bytecode, rendering conventional offset-based source mapping techniques ineffective.
Additionally, zero-knowledge proof systems require generating witness values corresponding to each execution step. This requirement introduces a multi-layered mapping problem spanning Solidity source code, LLVM IR, zkEVM bytecode, and constraint-level representations. Existing source mapping techniques do not accommodate this complexity.
3.5. Scope and Applicability
This work focuses on zkSolc, the compiler used by zkSync’s Layer 2 network, as a case study. However, the challenges and preliminary solutions explored in this work are relevant to other zero-knowledge proof systems and Layer 2 platforms, including Linea, Starknet, and Polygon zkEVM. Generalizing the framework to support these environments remains an avenue for future research.
4. Methodology
This section presents the design of a preliminary source mapping framework for zkSolc. The framework aims to enable reliable traceability between Solidity source code, LLVM IR, and zkEVM bytecode, addressing the traceability challenges introduced by zero-knowledge compilation pipelines. Our approach is intended to facilitate debugging, security auditing, and performance analysis in zk-Rollup environments.
4.1. Framework Overview
The proposed source mapping framework is integrated into the zkSolc compiler at both the frontend and backend stages. It introduces two primary mapping layers: one from Solidity to LLVM IR and another from LLVM IR to zkEVM bytecode. The first layer links Solidity constructs to their intermediate representations, while the second layer tracks how backend transformations affect these representations. Metadata is collected at each stage and consolidated into a unified source map to support execution trace reconstruction and error analysis.
4.2. Mapping Algorithm and Data Structures
The mapping process is designed to track transformations at each compilation stage, preserving traceability despite the optimizations introduced by LLVM and zk-SNARK-specific rewrites.
Frontend Mapping.
The process begins at the Solidity-to-LLVM IR stage. The compiler traverses the Solidity abstract syntax tree (AST) and annotates each node with source location metadata. Each AST node is mapped to a triple $(s,l,f)$, where $s$ is the starting position in the Solidity file, $l$ is the length of the corresponding code segment, and $f$ is the source file index. These annotations are propagated through the IR generation phase, ensuring that each LLVM instruction retains a reference to its originating AST node.
Backend Mapping.
Mapping from LLVM IR to zkEVM bytecode follows a similar approach but must additionally account for transformations introduced by LLVM optimization passes and zero-knowledge-specific rewrites. As IR instructions undergo transformations such as inlining, loop unrolling, and arithmetic simplification, corresponding bytecode offsets are dynamically recorded. Control flow modifications are also tracked to ensure that execution semantics can be reconstructed.
The mappings are stored in a unified table structured as $(s,l,f,I,B)$, where $I$ represents the associated LLVM IR instruction and $B$ denotes the corresponding bytecode offset. Additional metadata fields, including jump types, modifier depths, and zk-SNARK-specific annotations, are used to capture circuit-level transformations.
4.3. Integration into zkSolc
The mapping framework is integrated across multiple stages of the zkSolc compilation pipeline. In the frontend, AST nodes are extended to include source location annotations. The code generator is modified to propagate these annotations to LLVM IR instructions. In the LLVM backend, instrumentation modules monitor instruction-level transformations and associate bytecode offsets with their corresponding IR instructions. A dedicated mapping generator module consolidates the collected metadata into a unified mapping structure.
To enhance usability, a runtime query API is introduced, allowing developers to retrieve source mappings dynamically. This API is designed to be compatible with existing debugging environments such as Remix and Hardhat, enabling runtime errors to be traced back to Solidity source code.
4.4. Illustrative Example
To illustrate the mapping process, consider the submitVote function from the ZKVoting contract:
1function submitVote(bytes memory zkProof) external {
2 require(verifyZKProof(zkProof), "Invalid proof");
3 hasVoted[msg.sender] = true;
4}
During compilation, the Solidity parser generates an AST in which each node is annotated with source location information. The require statement and the state update are captured as distinct AST nodes. These nodes are then lowered into LLVM IR instructions:
1%1 = call i1 @verifyZKProof(%bytes* %zkProof)
2br i1 %1, label %valid, label %invalid
3%2 = load i1, i1* @hasVoted[msg.sender]
4store i1 true, i1* @hasVoted[msg.sender]
At the bytecode level, the mapping framework records the corresponding offsets:
0x10 -> CALLDATASIZE
0x14 -> CALL @verifyZKProof
0x18 -> JUMPDEST (label %valid)
0x20 -> MSTORE @hasVoted
The mapping information is stored in a structured table:
Table 1. Mapping Examples in ZKVoting Contract
Source Statement LLVM IR Instruction Bytecode Offset ZK Metadata
require(...); call @verifyZKProof 0x14 Constraint 1
hasVoted[msg.sender] = true; store i1 true 0x20 None
This mapping enables developers and auditors to correlate runtime execution behavior with Solidity source code and intermediate representations.
4.5. Preliminary Consistency Validation
To improve the reliability of the generated mappings, we incorporate a lightweight consistency validation process. This process includes syntactic and structural validation but does not aim to provide formal verification guarantees.
Syntactic Validation.
Syntactic validation ensures that the mapping structure adheres to predefined constraints. Specifically, the mapping table is defined as:
(1) $M=\{(s_{i},l_{i},f_{i},I_{i},B_{i})\mid 1\leq i\leq n\}$
where $s_{i}$ denotes the starting position of a Solidity segment, $l_{i}$ its length, $f_{i}$ the source file index, $I_{i}$ the corresponding LLVM IR instruction, and $B_{i}$ the associated bytecode offset. The validation process ensures that no two mappings overlap:
(2) $\forall i,j\quad(s_{i},l_{i})\cap(s_{j},l_{j})=\emptyset\quad\text{for }i\neq j.$
This condition avoids ambiguity in tracing execution, ensuring that each bytecode instruction corresponds to a distinct source-level segment.
Structural Validation.
Structural validation involves manually inspecting representative mappings to confirm that they correspond to meaningful Solidity constructs. During evaluation, we performed manual alignment checks between bytecode offsets and source code locations using debugging tools such as Remix IDE. While this process is not exhaustive, it provides a preliminary assessment of mapping correctness and highlights cases where transformations disrupt traceability.
Limitations.
We emphasize that the validation process described here is preliminary and informal. As discussed in Section 6, establishing semantic equivalence between Solidity constructs and their compiled representations is an undecidable problem in general. Therefore, our current validation focuses on structural consistency rather than formal verification.
5. Experiments
This section presents a preliminary evaluation of the proposed source mapping framework for zkSolc. Our evaluation focuses on two aspects: (i) mapping accuracy and (ii) compilation performance overhead, measured across both benchmark and real-world Solidity contracts.
5.1. Experimental Setup
Experiments were conducted on a workstation equipped with an Intel Core i9-12900K processor (16 cores, 3.9 GHz), 32 GB DDR5 RAM, running Ubuntu 22.04 LTS. We extended the publicly available zkSolc compiler (version 1.3.13) to include source mapping support. The modified compiler was compared against the unmodified zkSolc to assess compilation overhead and evaluate mapping correctness.
5.2. Evaluation Datasets
We used two datasets for evaluation:
  • Benchmark Dataset: A set of 50 manually selected Solidity contracts designed to cover a variety of language features, including loops, function overloading, inheritance, modifiers, and inline assembly. These contracts were chosen to represent common Solidity constructs.
  • Real-World Dataset: A collection of 500 deployed Solidity contracts obtained from the Ethereum mainnet and the zkSync Layer 2 network. These contracts span multiple application domains, including decentralized finance (DeFi), non-fungible tokens (NFTs), governance, and utility services. Contract sizes range from 50 to 3,000 lines of code.
5.3. Evaluation Methodology
We evaluated two key metrics:
Mapping Accuracy.
We define mapping accuracy as the percentage of bytecode instructions whose recorded source locations in the mapping table correctly match their corresponding Solidity statements. A mapping entry is considered correct if it satisfies the following criteria:
  1. (1)
    The recorded source location corresponds to the Solidity statement responsible for the execution of the bytecode instruction.
  2. (2)
    The mapping remains consistent with the control flow of the original source code.
To establish a ground truth for validation, we manually inspected a representative subset of the compiled contracts. For benchmark contracts, we cross-referenced mappings using debugging tools such as Remix IDE. For real-world contracts, we additionally compared our framework’s output against source maps generated by solc, where applicable.
Performance Overhead.
We measured the total compilation time for each contract with and without source mapping enabled. The performance overhead is reported as the percentage increase in compilation time introduced by the mapping framework.
5.4. Results
Table 2 provides an overview of the benchmark dataset, categorizing the contracts by feature type, mapping accuracy, and compilation performance. Table 3 extends this evaluation to real-world deployed contracts, offering insights into performance overhead across diverse application domains.
Table 2. Evaluation results for 50 benchmark Solidity contracts categorized by feature type. Compilation times are measured with and without source mapping enabled. Performance overhead refers to the percentage increase in compilation time.
Contract Type # Ctr. Avg. LOC Accuracy (%) Compilation Time (s) Perf. Overhead (%) No Mapping With Mapping Loops 6 461 97.32 1.85 2.31 10.67 Function Overloading 5 615 96.80 2.10 2.84 12.42 Inheritance 7 945 95.62 3.15 3.80 13.00 Modifiers 5 523 97.14 2.42 2.98 10.82 Inline Assembly 5 889 94.85 3.92 4.72 12.63 Event Logging 5 730 98.10 2.61 3.20 11.42 Storage Optimization 6 1115 96.42 3.75 4.50 13.57 Reentrancy Protection 5 860 97.65 3.35 4.20 11.94 Complex Control Flow 6 1196 94.90 4.01 5.00 12.99 Miscellaneous 5 682 96.88 3.05 3.87 11.90 Total / Average 50 842 96.47 3.03 3.84 11.92
Table 3. Evaluation results for 500 real-world Solidity contracts categorized by application domain. Compilation times are measured with and without source mapping enabled. Performance overhead refers to the percentage increase in compilation time.
Contract Type # Ctr. Avg. LOC Accuracy (%) Compilation Time (s) Perf. Overhead (%) No Mapping With Mapping DeFi (Lending, AMM, Staking) 71 1115 96.65 4.83 5.74 13.85 NFT Marketplaces / ERC Tokens 33 755 97.70 4.92 5.80 12.47 DAOs / Voting 58 1575 96.09 5.09 5.42 12.61 Utility (Multisig, Escrow) 47 1316 98.75 5.30 6.05 12.14 Gaming / Metaverse 55 2643 95.82 6.43 7.05 10.13 Identity / Privacy (KYC, zk-ID) 45 1422 98.42 4.35 5.12 11.67 Cross-Chain Bridges 39 1928 97.03 4.80 5.87 11.89 Enterprise Blockchain 31 2735 96.27 5.11 6.30 10.74 High-Frequency Trading 36 1276 98.86 4.25 5.08 12.95 Miscellaneous 35 1668 97.51 3.92 4.95 11.74 Total / Average 500 1573 97.20 4.81 5.94 11.92
The framework achieved an average mapping accuracy of 96.47% on the benchmark dataset and 97.20% on the real-world dataset. In both datasets, contracts with complex control flow (e.g., deep inheritance hierarchies and nested modifiers) and inline assembly exhibited slightly lower accuracy, reflecting the difficulty of maintaining traceability under aggressive compiler optimizations.
The compilation time overhead introduced by source mapping was moderate, averaging 11.92% across both datasets. Detailed measurement revealed that the Solidity-to-LLVM IR translation stage accounted for approximately 70% of the additional compilation time, while LLVM backend instrumentation and bytecode mapping contributed the remainder.
5.5. Alignment Criteria and Example
During validation, we considered a mapping entry to be correctly aligned if the recorded source location referred to the Solidity statement responsible for the corresponding bytecode instruction. For example, in the ZKVoting contract, the following mapping entries were considered correctly aligned:
Table 4. Alignment Validation Examples in ZKVoting Contract
Bytecode
Offset
Source Statement Alignment
0x14 require(verifyZKProof()) Aligned
0x20 hasVoted[msg.sender] = true Aligned
In cases where source constructs were eliminated or transformed during LLVM optimizations (e.g., constant folding or dead code elimination), no mapping entry was recorded.
5.6. Discussion and Key Insights
The evaluation results suggest that the proposed framework effectively captures source mappings in zk-Rollup compilation pipelines. The observed accuracy exceeded 96% in both benchmark and real-world datasets, indicating that source-level traceability is feasible despite the transformations introduced by zkSolc.
However, the results also highlight limitations. Contracts featuring inline assembly, complex modifiers, and unconventional control flow patterns exhibited slightly lower mapping accuracy, underscoring the difficulty of maintaining traceability in these scenarios. The compilation time overhead introduced by the mapping process remained within acceptable bounds, averaging approximately 12%.
We acknowledge that this evaluation remains preliminary. Mapping correctness was validated through manual inspection and comparison with solc mappings where possible. A fully automated and comprehensive validation framework is left for future work. Further analysis on larger datasets and more complex contracts will be necessary to generalize these findings.
6. Limitations
While the proposed source mapping framework for zkSolc demonstrates promising preliminary results, several limitations remain. These limitations highlight areas for further refinement and inform directions for future work.
One primary limitation is the compilation overhead introduced by the mapping process. Although our evaluation shows that the overhead remains moderate and acceptable for most development scenarios, it may pose challenges in high-throughput environments or automated deployment workflows involving frequent contract compilation. Further optimization of the mapping algorithm, particularly during the Solidity-to-LLVM IR translation phase, may help mitigate this overhead.
The current framework provides reliable traceability for standard Solidity constructs but exhibits reduced mapping accuracy when handling contracts with complex language features. Specifically, contracts involving deeply nested function calls, inline assembly, unconventional control flow, or experimental Solidity features present challenges for accurate traceability. Extending support for these advanced constructs without incurring significant performance penalties remains an open research problem.
Another practical limitation concerns integration with existing developer toolchains. Although the framework is designed to support execution trace reconstruction, it has not yet been fully integrated with widely used environments such as Remix, Hardhat, or Truffle. This limits its accessibility for developers accustomed to these platforms. Developing dedicated plugins or extensions to facilitate integration would substantially improve usability.
Furthermore, the current evaluation focuses exclusively on zkSolc and zk-Rollup-based compilation pipelines. The applicability of the framework to other zero-knowledge proof systems—such as Linea, Starknet, or Polygon zkEVM—has not been explored. While the underlying techniques may generalize to these platforms, confirming this requires further investigation and adaptation to system-specific compilation pipelines.
The framework’s mapping accuracy may also degrade in scenarios involving complex cross-contract interactions or deeply nested modifiers. Our evaluation identified minor inaccuracies in such cases, primarily due to compiler optimizations and non-linear control flow transformations. Improving mapping robustness under these conditions is an avenue for future enhancement.
Additionally, the current implementation focuses exclusively on static analysis and compilation-time mapping. It does not support runtime debugging features such as dynamic execution tracing or real-time error reporting. Incorporating such capabilities could significantly improve developer experience and is a promising direction for future work.
We also acknowledge that the evaluation datasets, although representative, may not fully capture all edge cases encountered in production environments. Expanding the evaluation to include a broader variety of contract types, code patterns, and execution scenarios would improve confidence in the framework’s generalizability.
Finally, the framework is tightly coupled with the current zkSolc architecture. Substantial changes to the compiler’s optimization strategies, IR transformations, or code generation processes may require corresponding modifications to the mapping framework to maintain accuracy and consistency.
Addressing these limitations will be the focus of future work. In particular, we plan to enhance mapping precision, reduce performance overhead, improve toolchain integration, and investigate generalization to other zero-knowledge compilation pipelines and execution environments.
7. Related Work
This section reviews prior research related to source mapping, smart contract debugging, and zero-knowledge Rollup compilation pipelines. While these areas have been studied independently, source mapping challenges specific to zero-knowledge-oriented compilers such as zkSolc remain underexplored.
7.1. Source Mapping and Debugging in Solidity
In Ethereum smart contract development, source mapping is a standard feature supported by mainstream Solidity compilers. For example, solc generates source maps that link compiled EVM bytecode instructions to corresponding Solidity source locations. This capability is fundamental to development environments and debugging tools such as Remix, Hardhat, and Truffle (Project, 2025b; Suite, 2025). Source mapping also plays a key role in formal verification workflows, enabling symbolic execution, execution trace reconstruction, and property checking.
Several analysis frameworks have been proposed to improve smart contract reliability and security. Static analyzers such as Slither (of Bits, 2025c), SmartCheck (SmartDec, 2025), and MythX (ConsenSys, 2025) detect common vulnerabilities at the source code level. Dynamic analysis tools, including Manticore (of Bits, 2025b) and Echidna (of Bits, 2025a), employ fuzzing and symbolic execution to expose runtime vulnerabilities. These tools fundamentally rely on accurate source mappings to correlate low-level execution behavior with high-level Solidity constructs. In compilation pipelines lacking source mapping support, such as zkSolc, debugging and auditing workflows are significantly hindered.
7.2. Zero-Knowledge Rollups and Compilation Tooling
Zero-knowledge Rollup (zk-Rollup) technologies have emerged as a leading solution for blockchain scalability (Ethereum Foundation, 2024). zk-Rollup systems, including zkSync (Labs, 2025b), process transactions off-chain and submit succinct cryptographic proofs to Ethereum for verification. Specialized compilers such as zkSolc have been developed to produce zero-knowledge-compatible bytecode tailored for these systems. Prior research in this space has predominantly focused on protocol-level efficiency, proof generation, and cryptographic correctness (Hacken, 2023). However, compiler-level challenges—particularly those concerning developer tooling and source-level traceability—have received comparatively limited attention.
A distinguishing characteristic of zero-knowledge compilers is their reliance on intermediate representations such as LLVM IR to facilitate aggressive optimization and proof-friendly transformations. These transformations introduce non-linear changes to control and data flow, complicating traceability between source code and compiled artifacts.
7.3. Source Mapping in Intermediate Representations
Source mapping between high-level languages and intermediate representations has been extensively studied in traditional compiler pipelines. For example, the LLVM framework (Lattner and Adve, 2004) supports debug metadata and source location annotations to enable traceability during optimization and code generation (Project, 2025a). Several studies have proposed techniques to preserve source mapping fidelity in the presence of conventional compiler optimizations, such as inlining and loop unrolling.
However, these techniques are not directly applicable to zero-knowledge compilation pipelines. The introduction of circuit-specific rewrites, arithmetic simplification at the bit level, and constraint generation for zk-SNARK proofs introduces additional layers of complexity beyond those encountered in general-purpose compilation.
7.4. Positioning and Contribution
Recent discussions within the zero-knowledge proof and smart contract development communities have emphasized the lack of robust debugging and traceability support in zk-Rollup ecosystems. To the best of our knowledge, no prior work has systematically addressed the problem of source mapping in the context of zero-knowledge-specific compilers such as zkSolc. Existing source mapping techniques in EVM or LLVM-based pipelines do not account for the non-linear transformations and circuit-level constraints inherent to zero-knowledge proof systems.
This paper takes an initial step toward addressing this gap. Our framework introduces a mapping mechanism tailored to the multi-stage compilation pipeline of zkSolc, providing preliminary support for tracing execution behavior across Solidity source code, LLVM IR, and zkEVM bytecode. While the approach does not offer formal correctness guarantees, it incorporates lightweight syntactic and structural validation to improve mapping consistency. We position this work as a foundation for future research toward more robust debugging, auditing, and verification toolchains in zero-knowledge smart contract development.
8. Conclusion
This paper presented a preliminary investigation into source mapping support for zkSolc, the compiler used to generate zero-knowledge-compatible smart contract bytecode. We proposed a source mapping framework that establishes traceability between Solidity source code, LLVM IR, and zkEVM bytecode, addressing a critical tooling gap in the zk-Rollup compilation pipeline.
Our evaluation on both benchmark and real-world Solidity contracts demonstrated that the proposed framework achieves high mapping accuracy while introducing only moderate compilation overhead. These results suggest that, despite the non-linear transformations inherent in zero-knowledge compilation workflows, meaningful source-level traceability can be recovered to support debugging, auditing, and performance analysis.
We acknowledge that this work constitutes an initial step toward robust source mapping in zero-knowledge environments. Several limitations remain, including reduced mapping precision in contracts featuring complex control flow and inline assembly, as well as integration challenges with existing developer toolchains. Furthermore, while we introduced lightweight syntactic and structural validation, formal guarantees of mapping correctness remain an open problem.
Future work will focus on addressing these limitations. Specifically, we plan to refine the mapping algorithm to improve accuracy for advanced Solidity constructs, optimize compilation overhead, and integrate the framework with popular development environments such as Remix and Hardhat. Extending the approach to other zero-knowledge compilation pipelines—such as those used by Linea, Starknet, and Polygon zkEVM—also represents an important direction for future research.
We hope that this preliminary investigation will lay the groundwork for further research on debugging, verification, and developer tooling in zero-knowledge smart contract ecosystems, ultimately contributing to a more transparent and reliable Layer 2 development experience.